# Stochastic Multigrid Minimization for Ptychographic Phase Retrieval

### v3

@@ Abstract
We propose a novel stochastic multigrid minimization method for ptychographic phase retrieval. In our formulation, the challenging nonconvex and ill-posed inverse problem is recast as the iterative minimization of a quadratic surrogate model that majorizes the original objective function. Our general framework encompasses the Ptychographic Iterative Engine (PIE) family of algorithms. By efficiently solving the surrogate problem using a multigrid method, our approach delivers significant improvements in both convergence speed and reconstruction quality compared with conventional PIE techniques.
@@ 1 Introduction
Mathematically, ptychographic phase retrieval is formulated as an inverse problem that is inherently nonconvex and ill-posed. To address these challenges, numerous algorithms have been proposed, including the well-known Ptychographical Iterative Engine (PIE) family [9, 10, 11], among others. In this work, we present a unifying framework that subsumes the entire PIE algorithm family and studies its convergence behavior, yielding practical guidance for devising more efficient solution strategies.
To further accelerate convergence, one promising strategy is to exploit multigrid methods. These methods, which have a long history in the numerical solution of partial differential equations, exploit the inherent multi-scale structure of the underlying problem by solving it on a hierarchy of discretizations [12]. Coarse-grid approximations capture the global structure of the solution, while fine-grid corrections resolve detailed features.
Among various general multigrid frameworks, the multilevel [^1] Throughout this work, the terms ”grid” and ”level” are used interchangeably. optimization algorithm (termed MG/OPT[13] has demonstrated significant improvements in convergence performance, while maintaining flexibility over a wide range of applications [14, 15, 16]. Specific to inverse problems, this hierarchical structure also helps mitigate the ill-posedness inherent in many inverse problems [17, 18, 19]. Recent work has adapted MG/OPT to the ptychographic context [20], resulting in reduced computational costs and enhanced convergence performance. Our proposed method extends the approach of [20]. We begin by recasting the problem as the iterative minimization of a quadratic surrogate that majorizes the original objective. Unlike coarse-grid models in general multigrid-based optimization, our surrogate is defined to be consistent with the fine-grid formulation, thereby eliminating the correction term typically required by MG/OPT. By incorporating stochastic sampling into this framework, we substantially reduce the cost of full-gradient evaluations and accelerate computational performance. Furthermore, our formulation seamlessly integrates with the state-of-the-art PIE family of ptychographic algorithms, enabling straightforward adaptation and implementation. Finally, our multigrid scheme supports automatic hyperparameter selection: once parameters are tuned on the finest grid, those for coarser levels are derived automatically, simplifying deployment and improving usability.
In summary, we propose an efficient stochastic multigrid minimization strategy for ptychographic phase retrieval that leverages a surrogate model to reformulate the inverse problem as a tractable stochastic optimization task, thereby significantly accelerating convergence. The remainder of this paper is organized as follows. Section 2 presents the mathematical foundation of ptychographic phase retrieval; Section 3 introduces the surrogate model and analyzes its convergence; Section 4 details the multigrid method and examines its convergence properties; and Section 5 presents numerical results.
@@ 2 Preliminaries
where $\theta(\cdot)$ denotes the element-wise principal argument [^2] Since $\theta(0)$ is undefined, at index $r$, we set $\left[\mathcal{R}_{k}(\bm{z}_{k})\right]_{r}=\left[\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\right)\right]_{r}$ when $\left[\mathcal{F}(\bm{Q}\odot\bm{z}_{k})\right]_{r}=0$..
@@
0:  initialization of object: $\bm{z}^{(0)}$, probe: $\bm{Q}$, measurements: $\{\bm{d}_{k}\}_{k=1}^{N}$, regularization constant: $\alpha$.
@@ 3 Surrogate Minimization
where $\mathcal{R}_{k}$ is defined in Definition 2.1. We note that this term is ill-defined if $\theta\left(\left[\mathcal{F}\left(\bm{Q}P_{k}\bm{z}^{(j)}\right)\right]_{r}\right)=0$ at some index $r$, so that we complement the original definition with the following specification,
$\left[\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right]_{r}=\begin{cases}\left[\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\right)\right]_{r}\,,&\text{if }j=0\,,\\ \left[\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j-1)}\right)\right]_{r}\,,&\text{if }j\geq 1\,.\end{cases}$ (8)
@@ 3.1 Properties of the surrogate model
Figure 2: Illustration of the majorization property.
@@
The proof for this proposition is included in Appendix B.1. The inequality (12) is the reason for the method to be termed majorization.
@@
The proof for this theorem is included in Appendix B.2. We note that the specification made in Eqn. (8) is necessary in the proof for the optimality condition.
@@
The proof for this theorem is included in Appendix C. We note that the strategy in [26] has played an instrumental role in the proof.
4 Stochastic Multigrid Surrogate Minimization
@@
4.1 Multilevel–Adaptive–Guided Ptychographic Iterative Engine (MAGPIE)
@@
0:  Initialization of object: $\bm{z}^{(0)}$, probe: $\bm{Q}$, measurements: $\{\bm{d}_{k}\}_{k=1}^{N}$, regularization term: $\bm{u}$.
@@ 4.2 Converging properties of MAGPIE and weight selection
$\bm{W}_{\bm{z}}=\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,,\quad\bm{W}_{\mathcal{R}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot|\bm{Q}|^{2}}{\bm{Q}\odot\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}\,,\quad\text{and}\quad\bm{W}_{\bm{u}}^{H}=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$ (16)
As will be shown more clearly in Propositions 4.1, 4.2, and 4.4, and in Theorem 4.3, these specific forms not only prevent singularities, maintain coarse-to-fine consistency, and preserve the descent property, but also ensure stability. It is worth noting that the quantity $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ appears in both $\bm{W}_{\bm{z}}$ and $\bm{W}_{\mathcal{R}}$. Its physical meaning is straightforward: in each $2\times 2$ bin of $|\bm{Q}|^{2}$, $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ computes the average of the four entries and upsamples that average back to the original four positions, whereas $|\bm{Q}|^{2}$ each entry’s original value. As a result,
@@
See Appendix E for the proof.
@@
The proof for this proposition is included in Appendix F.
@@
The proof for this theorem is included in Appendix G.1.
Whenever a model is represented on both fine and coarse grids, parameters fixed on the fine grid must have consistent counterparts on the coarse grid. We show below that, with $\bm{W}_{\bm{z}},\bm{W}_{\mathcal{R}},\bm{W}_{\bm{u}}^{H}$ designed as in (16), the regularization term can be transferred automatically between levels, requiring no manual tuning for the coarse-grid regularization term $\bm{u}_{H}$.
@@
The proof for this proposition is included in Appendix H.
@@ 5 Numerical Examples
We first describe the problem setup, and then in Sections 5.1, 5.2, and 5.3, we evaluate the performance of the algorithms from three perspectives: the number of multigrid levels, robustness to noise, and the effect of the overlap ratio. Finally, Section 5.4 evaluates the algorithm on a simulated integrated-circuit object to illustrate its practical imaging performance.
@@ 5.1 Effect of multigrid levels on performance
In this experiment, we investigate how the number of coarse levels impacts the performance of MAGPIE. In particular, with $l=1,\cdots,\log_{2}(m)$ levels, we name the corresponding method $\texttt{MAGPIE}\_l$. Since $m=128$, the maximum level for MAGPIE is $\log_{2}(m)=7$.
@@
The residuals and errors are presented in Figure 6 on the log-log scale. We note that $\texttt{MAGPIE}\_l$ outperforms rPIE and L-BFGS, and converges faster with each additional coarse level. Therefore, for subsequent tests, we consider only MAGPIE at its deepest level, $l=\log(m)$.
@@ 7 Acknowledgments
$\nabla_{\bm{z}_{k}^{\prime}}\Phi^{q,r}_{\bm{z}_{k}}(\bm{z}_{k}^{\prime})=\overline{\bm{Q}}\odot(\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k}))+\bm{u}\odot(\bm{z}_{k}^{\prime}-\bm{z}_{k})\,,$
@@
$\Phi^{q,r}_{\bm{z}_{k}}(\bm{z}_{k}^{\prime})=\frac{1}{2}\left\|\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}+\frac{1}{2}\bm{u}\cdot|\bm{z}_{k}^{\prime}-\bm{z}_{k}|^{2}$
@@
$\frac{\partial\,\left\|\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}}{\partial\,\bm{z}_{k}^{\prime}}$ $=\frac{\partial\,\left(\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right)\cdot\overline{\left(\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right)}}{\partial\,\bm{z}_{k}^{\prime}}$
$=\overline{\left(\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right)}\operatorname{diag}(\bm{Q})\,.$
@@
$\frac{\partial\,\bm{u}\cdot|\bm{z}_{k}^{\prime}-\bm{z}_{k}|^{2}}{\partial\,\bm{z}_{k}^{\prime}}$ $=\frac{\partial\,(\bm{z}_{k}^{\prime}-\bm{z}_{k})\cdot\left(\bm{u}\odot\overline{(\bm{z}_{k}^{\prime}-\bm{z}_{k})}\right)}{\partial\,\bm{z}_{k}^{\prime}}$
$=\bm{u}\odot\overline{(\bm{z}_{k}^{\prime}-\bm{z}_{k})}\,.$
@@
$\nabla_{\bm{z}_{k}}\Phi^{q,r}_{\bm{z}_{k}}$ $=2\frac{\partial\,\Phi^{q,r}_{\bm{z}_{k}}}{\partial\,\bm{z}_{k}}^{*}\,,$
$=2\cdot\frac{1}{2}\cdot\left(\operatorname{diag}\left(\overline{\bm{Q}}\right)\left(\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right)+\bm{u}\odot(\bm{z}_{k}^{\prime}-\bm{z}_{k})\right)\,,$
$=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}^{\prime}-\mathcal{R}_{k}(\bm{z}_{k})\right)+\bm{u}\odot(\bm{z}_{k}^{\prime}-\bm{z}_{k})\,.$
@@
    Since at the zero entries $\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)_{m}$, we impose $\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)=\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)$, we have, for all $k=1,2,\dots,N$,

### v4

# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

@@ Abstract
We introduce MAGPIE (Multilevel––Adaptive––Guided Ptychographic Iterative Engine), a stochastic multigrid solver for the ptychographic phase-retrieval problem. The ptychographic phase-retrieval problem is inherently nonconvex and ill-posed. To address these challenges, we reformulate the original nonlinear and nonconvex inverse problem as the iterative minimization of a quadratic surrogate model that majorizes the original objective. This surrogate not only ensures favorable convergence properties but also generalizes the Ptychographic Iterative Engine (PIE) family of algorithms. By solving the surrogate model using a multigrid method, MAGPIE achieves substantial gains in convergence speed and reconstruction quality over traditional approaches.
@@
Mathematically, ptychographic phase retrieval is formulated as an inverse problem that is inherently nonconvex and ill-posed. To address these challenges, numerous algorithms have been proposed, including the well-known Ptychographical Iterative Engine (PIE) family [9, 10, 11]. This family of algorithms casts the problem in an iterative fashion and stabilizes the reconstruction through adding a quadratic regularization term.
Another feature of the ptychographic phase retrieval problem is its inherent multiscale structure. Coarse and fine features coexist in the to-be-reconstructed object, requiring different levels of numerical accuracy. This naturally calls for a multigrid solver. Multigrid methods have a long history in the numerical solution of partial differential equations, exploiting the inherent multi-scale structure of the underlying problem by solving it on a hierarchy of discretizations [12]. Coarse-grid approximations capture the global structure of the solution, while fine-grid corrections resolve detailed features. In the context of optimization, a multilevel [^1] Throughout this work, the terms “grid” and “level” are used interchangeably. optimization algorithm (termed MG/OPT) [13] has demonstrated significant improvements in convergence performance, while maintaining flexibility over a wide range of applications [14, 15, 16]. Specific to inverse problems, this hierarchical structure also helps mitigate the ill-posedness inherent in many inverse problems [17, 18, 19]. Recent work has adapted MG/OPT to the ptychographic context [20], resulting in reduced computational costs and enhanced convergence performance.
In this paper, we propose MAGPIE (Multilevel–Adaptive–Guided Ptychographic Iterative Engine), a solver that integrates the advantages of both PIE and MG/OPT. As done in PIE, we recast the problem into an iterative minimization problem with each step solving a quadratic surrogate that majorizes the original objective. The majorization effect brings us a favorable convergence rate. Meanwhile, we also deploy a multigrid-based optimization strategy and reconstruct the coarse and fine features of the object level by level. Unlike other multigrid-based optimization methods whose coarse-level models are generically incompatible with those on the fine level, our proposed surrogate naturally builds compatibility across scales, thereby eliminating the correction term typically required by MG/OPT. This compatibility also ensures automatic hyperparameter selection: once parameters are tuned on the finest grid, those for coarser levels are derived automatically, simplifying tuning in the implementation and improving usability. The overall method is run in a stochastic fashion, and the final algorithm delivers nice reconstruction with substantially reduced cost of full-gradient evaluations and accelerated computational performance.
The remainder of this paper is organized as follows. Section 1 presents the ptychographic phase retrieval problem and introduces two main techniques to be deployed: the surrogate modeling deployed in the PIE family algorithm and multigrid optimization strategy; Section 2 introduces our surrogate model, a model that will be shown to majorize the original objective and thus brings in favorable convergence properties; in Section 3 we finally present our ultimate solver: MAGPIE, through designing a multigrid method for the surrogate model. We also examine the iterative performance of the algorithm. In Section 4 we present numerical results.
@@ 1 Preliminaries
where $\theta(\cdot)$ denotes the element-wise principal argument [^2] Since $\theta(0)$ is undefined, at index $r$, we set $\left[\theta\left(\mathcal{F}(\bm{Q}\odot\bm{z}_{k})\right)\right]_{r}=0$ when $\left[\mathcal{F}(\bm{Q}\odot\bm{z}_{k})\right]_{r}=0$..
@@
0:  Object: $\bm{z}^{(0)}$, probe: $\bm{Q}$, measurements: $\{\bm{d}_{k}\}_{k=1}^{N}$, regularization constant: $\alpha$.
@@ 2 Surrogate Minimization
where $\mathcal{R}_{k}$ is defined in Definition 1.1. We note that in the definition of $\mathcal{R}_{k}$, the term $\left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}$ is ill-defined if $\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0$ at some index $r$, so that we complement the original definition with the following specification:
$\left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}=\begin{cases}0\,,&\text{if }j=0\,\&\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j-1)}\right)\right)\right]_{r}\,,&\text{if }j\geq 1\,\&\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}\,,&else\end{cases}$ (8)
@@ 2.1 Properties of the surrogate model
Figure 2: Illustration of the majorization property adapted from [26].
@@
The proof of this theorem appears in Appendix B.1. The first statement is straightforward to verify. To show the inequality (12), we deploy the definitions in Eqn. (1) and Eqn. (7) and prove the inequality for every $\Phi_{k}$. The last statement comes from straightforward computation. We call $\widetilde{\Phi}$ a majorization of $\Phi$ due to the inequality Eqn. (12).
@@
We leave the proof of this theorem to Appendix B.2. The monotonicity inequality follows directly from Proposition 2.1, and the first-order optimality of $\Phi$ follows from the corresponding first-order optimality condition on the quadratic surrogate, with $\mathcal{R}_{k}$ specified in (8)
@@
We leave the proof of this theorem to Appendix C. The overall derivation follows that in [27], where the increments at each iteration are bounded, and a telescoping-sum argument is applied.
3 MAGPIE
@@
3.1 Solver development of MAGPIE
@@
0:  Object: $\bm{z}^{(0)}$, probe: $\bm{Q}$, measurements: $\{\bm{d}_{k}\}_{k=1}^{N}$, regularization term: $\bm{u}$.
@@ 3.2 Converging properties of MAGPIE and weight selection
$\bm{W}_{\bm{z}}=\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,,\quad\bm{W}_{\mathcal{R}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot|\bm{Q}|^{2}}{\bm{Q}\odot\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}\,,\quad\bm{W}_{\bm{u}}^{H}=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$ (16)
As will be shown more clearly in Propositions 3.1, 3.2, and 3.4, and in Theorem 3.3, these specific forms not only prevent singularities, maintain coarse-to-fine consistency, and preserve the descent property, but also ensure stability. It is worth noting that the quantity $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ appears in both $\bm{W}_{\bm{z}}$ and $\bm{W}_{\mathcal{R}}$. Its physical meaning is straightforward: in each $2\times 2$ bin of $|\bm{Q}|^{2}$, $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ computes the average of the four entries and upsamples that average back to the original four positions, whereas $|\bm{Q}|^{2}$ each entry’s original value. As a result,
@@
To prove this proposition, we realize that $\mathbf{W}_{\mathbf{z}}$ represents the weighted average within a bin and the bin size is $4$. The control for the other two terms comes from Jensen’s inequality used on quadratic functions. The proof is left to Appendix E.
@@
We leave the proof to Appendix F. Most of the proof is a straightforward derivation.
@@
The proof of this theorem appears in Appendix G.1. First, we state Proposition G.1 showing a specific form under which the coarse-grid update is a descent direction. Then we show that the coarse-grid update in Eqn. (17) indeed falls into that form.
Whenever a model is represented on both fine and coarse grids, parameters fixed on the fine grid must have consistent counterparts on the coarse grid. We show below that, with $\bm{W}_{\bm{z}},\bm{W}_{\mathcal{R}},\bm{W}_{\bm{u}}^{H}$ designed as in Eqn. (16), the regularization term can be transferred automatically between levels, requiring no manual tuning for the coarse-grid regularization term $\bm{u}_{H}$.
@@
The proof of this proposition appears in Appendix H. The key step uses Eqn. (17), which links the coarse-grid update to the fine-grid surrogate gradient.
@@ 4 Numerical Examples
We first describe the problem setup, and then in Sections 4.1, 4.2, and 4.3, we evaluate the performance of the algorithms from three perspectives: the number of multigrid levels, robustness to noise, and the effect of the overlap ratio. Finally, Section 4.4 evaluates the algorithm on a simulated integrated-circuit object to illustrate its practical imaging performance.
@@ 4.1 Effect of multigrid levels on performance
In this experiment, we investigate how the number of coarse levels impacts the performance of MAGPIE. In particular, with $l=1,\cdots,\log_{2}(m)$ levels, we name the corresponding method MAGPIE_$\_l$. Since $m=128$, the maximum level for MAGPIE is $\log_{2}(m)=7$.
@@
The residuals and errors are presented in Figure 6 on the log-log scale. We note that MAGPIE_$\_l$ outperforms rPIE and L-BFGS, and converges faster with each additional coarse level. Therefore, for subsequent tests, we consider only MAGPIE at its deepest level, $l=\log(m)$.
@@ 6 Acknowledgments
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,,$
@@
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$
@@
$\frac{\partial\,\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}}{\partial\,\bm{z}_{k}}$ $=\frac{\partial\,\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\cdot\overline{\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)}}{\partial\,\bm{z}_{k}}$
$=\overline{\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)}\operatorname{diag}(\bm{Q})\,.$
@@
$\frac{\partial\,\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}}{\partial\,\bm{z}_{k}}$ $=\frac{\partial\,\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\cdot\left(\bm{u}\odot\overline{\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)}\right)}{\partial\,\bm{z}_{k}}$
$=\bm{u}\odot\overline{\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)}\,.$
@@
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}$ $=2\frac{\partial\,\Phi^{\texttt{PIE}}_{k}}{\partial\,\bm{z}_{k}}^{*}\,,$
$=2\cdot\frac{1}{2}\cdot\left(\operatorname{diag}\left(\overline{\bm{Q}}\right)\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\right)\,,$
$=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,.$
@@
    Since at the zero entries $\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)$, we impose $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ in the Eqn. (8), we have, for all $k=1,2,\dots,N$,